A ball of mass 50g is thrown vertically upwards with a velocity of 20m/s and later caught by the thrower.

(a) Draw a velocity – time graph for the motion of the ball indicating the values of time.

(b) Determine the total distance traveled during its journey?



1
Expert's answer
2021-03-04T11:51:10-0500

a)



b)


"d=2h\\v^2=2gh\\
d=\frac{v^2}{g}\\d=\frac{20^2}{10}=40\ m"